# Dual-Level Boost Network for Long-Tail Prohibited Items Detection in X-ray Security Inspection

### v1

@@ 4.4.2 Impact of Hyperparameters
To evaluate the sensitivity of the augmentation strength parameter $k$ on model-level augmentation, we experimented with $k$ values of 2, 4, 8, and 16, as shown in Figure 4. $k$ equals 2 achieved an AP50 of 81.3%. The lower value of $k$ limits the amount of contextual information available for the augmentation, which results in less comprehensive relational learning. $k$ equals 4 yielded the highest AP50 of 81.8%, demonstrating that moderate incorporation of contextual features strikes the best balance between providing meaningful relationships without overwhelming the model with noise. $k$ equals 8 resulted in a slight decline to AP50 of 81.6%. Here, the increase in contextual information started to introduce noise, which diluted the effectiveness of the relational augmentation, indicating that too much contextual data can make feature learning less focused. $k$ equals 16 further decreased the performance to AP50 of 81.0%, with performance for tail classes like Class NL dropping to 14.6%. This suggests that excessive augmentation leads to model confusion, especially for tail classes in real-world scenario, as it becomes harder for the model to distinguish between relevant and irrelevant contextual cues of surroundings.
@@ 5 Conclusion
In conclusion, the proposed DBNet addresses critical challenges in X-ray prohibited item detection, particularly the issues stemming from the long-tail distribution of item appearances and the inherent complexities of X-ray imagery. By leveraging a dual-level boost framework, combining data augmentation via Poisson blending and a context-aware feature enhancement module, DBNet significantly improves the detection of rare, infrequent items while maintaining high accuracy for common ones. Extensive experiments demonstrate that DBNet outperforms state-of-the-art approaches, especially in enhancing detection for underrepresented classes. The contribution of this paper marks a significant advancement towards automated and robust security inspection systems, ensuring enhanced public safety in real-world environments like public transportation hubs.

### v2

@@ 4.6.2 Impact of Hyperparameters
To evaluate the sensitivity of the augmentation strength parameter $k$ on model-level augmentation, we experimented with $k$ values of 2, 4, 8, and 16 on RetinaNet [26], as shown in Fig. 4. $k$ equals 2 achieved an AP50 of 81.3%. The lower value of $k$ limits the amount of contextual information available for the augmentation, which results in less comprehensive relational learning. $k$ equals 4 yielded the highest AP50 of 81.8%, demonstrating that moderate incorporation of contextual features strikes the best balance between providing meaningful relationships without overwhelming the model with noise. $k$ equals 8 resulted in a slight decline to AP50 of 81.6%. Here, the increase in contextual information started to introduce noise, which diluted the effectiveness of the relational augmentation, indicating that too much contextual data can make feature learning less focused. $k$ equals 16 further decreased the performance to AP50 of 81.0%, with performance for tail classes like Class NL dropping to 14.6%. This suggests that excessive augmentation leads to model confusion, especially for tail classes in real-world scenario, as it becomes harder for the model to distinguish between relevant and irrelevant contextual cues of surroundings.
@@
4.7 More Analysis and Discussion
This section discusses the scalability of manual segmentation and the generalization of to transformer-based models.
4.7.1 Manual segmentation scalability
In our initial experiments, we tested SAM [17] for automatic segmentation, but the performance was significantly lower than when manually annotating the objects. This discrepancy is particularly noticeable in terms of accuracy, as SAM [17] struggled with distinguishing some fine-grained details, leading to less precise segmentations (examples are in Fig. 5). However, the manual annotation of 78 instances, as shown in the supplementary material, took less than an hour, which is far less than the time required for a single training epoch. Given this, the manual segmentation process presents an efficient trade-off between annotation time and the quality of augmented data, making it a practical solution for the current scope of our work.
Figure 5: SAM automatic annotation (left: original, right: SAM).
4.7.2 Generalization on Transformer-based Models
We further evaluated the performance of our method by incorporating it with a transformer-based model, DINO [43]. As shown in Tab 5, our augmentation strategy leads to a noticeable improvement in detection performance. Specifically, the baseline DINO model achieves an AP50 of 84.4, whereas the integration of our technique boosts this to 86.3, with a significant increase in tail categories (NL) from 26.4 to 37.7. This demonstrates that our approach is effective not only for traditional convolutional models but also generalizes well to transformer-based models, improving overall performance, particularly for underrepresented categories.
Method AP50 AP50 across various categories
PO1 PO2 WA LA MP TA CO NL
DINO [43] 84.4 96.3 95.7 94.1 98.0 98.2 96.4 69.7 26.4
+ours 86.3 96.2 96.1 93.7 98.3 98.2 96.1 74.0 37.7
Table 5: Comparison with the transformer-based model, DINO.
@@ 5 Conclusion
In conclusion, the proposed XIDNet addresses critical challenges in X-ray prohibited item detection, particularly the issues stemming from the long-tail distribution of item appearances and the inherent complexities of X-ray imagery. By leveraging a dual-level boost framework, combining data augmentation via Poisson blending and a context-aware feature enhancement module, XIDNet significantly improves the detection of rare, infrequent items while maintaining high accuracy for common ones. Extensive experiments demonstrate that XIDNet outperforms state-of-the-art approaches, especially in enhancing detection for underrepresented classes. The contribution of this paper marks a significant advancement towards automated and robust security inspection systems, ensuring enhanced public safety in real-world environments like public transportation hubs.